OpenStudy (anonymous):

A 2.9×103 kg car accelerates from rest under the action of two forces. One is a forward force of 1153 N provided by traction between the wheels and the road. The other is a 920 N resistive force due to various frictional forces. How far must the car travel for its speed to reach 2.3 m/s? Answer in units of m

OpenStudy (jziggy):

If we define the x direction as positive in the same direction as the accelleration \[\sum_{}^{}Fx = 1153N - 920N = M*a\] \[233N = (2.9*103kg)*a\] \[233N = (298.7kg)*a\] \[233N/298.7kg = a\] \[a \approx 0.78005(m/s) \] Using the equation: \[v = vi + a*t\] \[2.3(m/s) = 0.78005(m/s^2)t\] \[t \approx 2.94853(s)\] Using the equation: \[d = vi*t + (t*a^2)/2\] \[d \approx 1.79412(m)\] Therefore, the car must travel approximately 1.79 meters to reach a speed of 2.3m/s

OpenStudy (anonymous):

I tried that answer and it didn't work :/

OpenStudy (anonymous):

I think jziggy is right. but I think there was a mix up from the post to the solution. the phrase: "A 2.9×103 kg car" should have been "A 2.9×10^3 kg car" I worked it out and it turns out: A 2.9×103 kg car = 658 Lb car [which is lighter than a motorcycle] A 2.9×10^3 kg car = 6,393 Lb car [which is more believable] So, from the second line of the equations posted, all the way to the end can be re-worked in the same way using the car mass of 2900 kg to get the right answer. ^_^

OpenStudy (anonymous):

I tried to rework the problem but I think I may have done something wrong O.o

OpenStudy (anonymous):

what do you get for a? and for t?

OpenStudy (anonymous):

@RavenLynette

OpenStudy (anonymous):

for t I got 28.6265 and for a I got .080345 :/

OpenStudy (anonymous):

l or t is 0.080345

OpenStudy (anonymous):

Maybe that's right

OpenStudy (anonymous):

time is that?

OpenStudy (anonymous):

a is right 0.0803

OpenStudy (anonymous):

t = 28.64 is correct

OpenStudy (anonymous):

I got .09 something for the distance

OpenStudy (anonymous):

distance should be = .5*.0803*28.6^2 = 28.67 meters \[d=vi∗t+(1/2)*a*t^{2}\]

OpenStudy (anonymous):

@RavenLynette you 'see' it?

OpenStudy (anonymous):

Oh ok thanks :)

OpenStudy (anonymous):

^_^